OpenStudy (anonymous):

A small portable TV in the store has dimensions of 12 inches by 13 inches. What is its diagonal measure? Between 15 and 16 cm Between 16 and 17 cm Between 17 and 18 cm

jhonyy9 (jhonyy9):

diagonal D=sqrt(144+169)=sqrt(313)= 17,69 - so the last will be right ,between 17 and 18 cm